### v1

@@ III.2 Compressed spectra
When constructing ROMs for our inhomogeneous Schrödinger equations, Eq. (9) turns into Eq. (18), with $M=E-H$. Per Eqs. (18) and (23), the resulting $\begin{bmatrix}\mathcal{A}(E)\end{bmatrix}$ has simple poles located at the zeros of $\det(\begin{bmatrix}M\end{bmatrix})$ in the complex $E$ plane. I.e., these pole positions are determined by the eigenvalues of the following generalized eigenvalue problem:
@@
When two conditions are satisfied: $|{S}\rangle=|\tilde{{S}}\rangle$ and $H$ being Hermitian in the training calculations, as held in computing elastic scattering and inclusive response functions, $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ become complex symmetric matrices. This property is the shared characteristic in the existing NHQM methods Moiseyev (2011); Michel and Płoszajczak (2021). The training equations turn into,
$(E_{\alpha}-H_{\alpha})|\Psi_{\alpha}\rangle$ $=|{S}_{\alpha}\rangle\ ,$ (31)
$(E_{\alpha}^{\ast}-H_{\alpha})|\tilde{\Psi}_{\alpha}\rangle$ $=|{S}_{\alpha}\rangle\ .$ (32)
Since $|\Psi\rangle$ is analytic in terms of $E$, we have
$|\tilde{\Psi}_{\alpha}\rangle=|\Psi(E_{\alpha}^{\ast},\bm{\theta}_{\alpha})\rangle=\left(|\Psi(E_{\alpha},\bm{\theta}_{\alpha})\rangle\right)^{\dagger}\ .$ (33)
Therefore, the inner product with complex conjugation in computing $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ turns into the so-called $C$-product (labeled with c subscript) Moiseyev (2011):
$\begin{bmatrix}N\end{bmatrix}_{\alpha\beta}$ $=\langle\Psi_{\alpha}|\Psi_{\beta}\rangle_{c}=\begin{bmatrix}N\end{bmatrix}_{\beta\alpha}\ ,$ (34)
$\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}$ $=\langle\Psi_{\alpha}|H|\Psi_{\beta}\rangle_{c}=\begin{bmatrix}H\end{bmatrix}_{\beta\alpha}\ .$ (35)
This proves the previous assertion. Note that $H$ is Hermitian in the training calculations, but $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ are non-Hermitian in general and complex symmetric in this particular case.
As briefly mentioned at the beginning of Sec. III and elaborated later in Secs. IV and V, these eigenvalues are distributed in a pattern similar to the pole distribution of a (near)-optimal rational approximation discussed in Sec. II.3. The emulation performance certainly depends on whether $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ can reproduce the physics, including the pole positions, in the region where the bound states and resonances are located.
However, the DC states in the emulator’s compressed spectrum depend on the setup of the training points, in particular, the distribution of $E^{tr}_{\alpha}$ in the complex $E$ plane. As one interesting specific case, when (1) ${S}=\tilde{{S}}$ and (2) both $E^{tr}_{\alpha}$ and $E^{tr\ast}_{\alpha}$ are included as the training points associated with a given $\bm{\theta}_{\alpha}$, the eigenenergies are always real and thus the discretized BCTs are along the real axis. See Fig. 3 for a visual presentation. A heuristic proof, making use of the so-called pseudo-Hermiticity concept Mostafazadeh (2002a) and Mostafazadeh (2002a), is provided in the Appendix A. Such a setup could be beneficial for discrete bound spectrum calculations. However, the discretized BCTs should be away from the real axis for the continuum calculations. Therefore, we focus on the emulations with $\mathrm{Im}\,E^{tr}_{\alpha}>0$ in the training points.
@@ III.3 Difference between the existing NHQM and the CEE methods
We are now at a good point to discuss the differences between the existing NHQM methods and our CEE. The previous approaches must analytically continue the conventional single-particle basis (e.g., harmonic oscillator basis) in which $H$ is Hermitian to a basis that $H$ turns non-Hermitian. These methods then construct a many-body basis as a direct product of the single-particle basis. The numerical difficulty could arise during the analytical continuation, limiting how far away from the real axis the BCTs can be moved.
In contrast, CEE achieves the “non-Hermitization” of $H$ by building the many-body basis from the solution snapshots of the inhomogeneous Schrödinger equations with complex energies, i.e., by the RBM-based subspace projection (or viewed as spectrum compression) during the on-line emulation stage. At the off-line training stage, we don’t need to—but we can—perform analytical continuation of the $H$ matrix to obtain the training solution snapshots, as the solutions to the equation with a Hermitian $H$ already have bound-state-like boundary conditions. As a result, the non-Hermitian $\begin{bmatrix}H\end{bmatrix}$ matrix is generally much lower-dimensional than the corresponding matrix in the other NHQM methods.
It is also relevant to note that our approach targets a specific portion of the spectrum, as the eigenenergies of the compressed spectrum are distributed in a finite domain of the complex $E$ plane. The portion is not precisely defined; however, when the $E$ variable is extrapolated too far away from the range of the training energies during the emulation stage, the emulation errors become significant. In contrast, in the existing NHQM methods, the spectrum range is defined by the energy cut-off applied on a single-particle basis.
@@ III.4 Some numerical details
When performing the $\mathcal{A}$ emulation, we need to solve Eq. (18) for $c$ and $\tilde{c}$. However, the equation could become ill-conditioned when $N_{b}$ is too large, which, in fact, reflects the efficiency of the dimension reduction. We apply the singular value decomposition (SVD) decomposition to the $\begin{bmatrix}M\end{bmatrix}$ and eliminate the singular values $\sigma_{.}$ that satisfy $|\sigma_{i}/\sigma_{max}|<10^{-13}$, a threshold similar to the relative errors of the training calculation, to perform pseudo-inverse to compute $c$ and $\tilde{c}$. When solving the generalized eigenvalue problem to emulate spectra using Eq. (30), we invert $\begin{bmatrix}N\end{bmatrix}$ using the pseudo-inverse with the same SVD singular value truncation.
Another way exists to emulate $\mathcal{A}$ in $E$ when we have an emulated spectrum at $\bm{\theta}$. We note that
@@
with $E_{\beta}$ as the emulated eigenvalues. To fix all the residues, $W_{\beta}$, we first compute $\mathcal{A}(E_{\alpha}^{tr},\bm{\theta})$ by using the first $\mathcal{A}$ emulation method. The residues can be computed by solving a linear equation system:
$\begin{bmatrix}\mathcal{A}(E_{\alpha}^{tr},\bm{\theta})\end{bmatrix}$ $=\sum_{\beta=1}^{N_{b}}\frac{W_{\beta}(\bm{\theta})}{E_{\alpha}^{tr}-E_{\beta}(\bm{\theta})}\ .$ (37)
@@ III.5 Existing works
In the MOR studies of dynamic linear systems, the equations in the frequency domain take the form of Eq. (9). The so-called rational-Krylov method (e.g., Chap. 10 and 11 in Ref. Antoulas (2005)) is similar to our CEE. However, these studies are about systems with finite-dimensional matrices. Similarly, recent two studies Van Beeumen et al. (2017); Peng et al. (2019) in quantum chemistry also apply this method to emulate $\mathcal{A}(E)$ involving continuum physics. Again, the critical difference between finite-dimensional matrix and $H$ with continuum was not studied there, including the point that the continuum spectra can be efficiently compressed. It seems the emulated eigenenergies in these works are always real, i.e., their Galerkin projection, used to derive the ROMs, differs significantly from the one used in this work. In addition, we demonstrate the effectiveness of emulation for other parameters, which has not been studied in Refs. Van Beeumen et al. (2017); Peng et al. (2019).
@@ IV Two-body demonstrations
In this section, we work with a toy model for two-nucleon-like systems. The particle mass is $M_{N}=940$ MeV in the natural units. The interaction, in the $\ell$th partial-wave channel, takes a separable form:
@@
The form factor in the momentum-space representation, $g_{\ell}(q)$, is a Gaussian function with a width parameter $\Lambda_{2}$; $\lambda$ is the coupling-strength parameter. Here, we focus on $s$ and $p$ waves. The associated $\mathcal{A}(E)$s are known analytically in the complex $E$ plane. The details about the model and relevant analytical formulas can be found in Appendix B.
For emulation, the training calculations are carried out using a bound-state method, employing the Lagrange function of the Legendre polynomials as basis Baye (2015). The basis function is defined in a finite interval of the spatial radial coordinate, $[0,R_{IR}]$, as mentioned in discussing Fig. 1. We choose $R_{IR}$ large enough and the UV resolution scale small enough so that the emulator components, $\begin{bmatrix}N\end{bmatrix}$ and $\begin{bmatrix}H\end{bmatrix}$, all converged to a level of $10^{-12}$ or even smaller.
@@ IV.1 Emulation with fixed $H$ but varying sources
Figure 2: Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{rel})$. The training wave functions are plotted in the top panel, and at the bottom, the emulated and the exact wave functions are compared for a particular point for on-shell scattering. The information about training points is listed in the top panel, while the information about the testing point is provided in the title of the bottom panel.
@@
In this subsection, we hope to gain some understanding of our emulators from the wave function perspective. We perform emulation in three-dimensional parameter space for $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{rel})$. $E_{rel}$ is a parameter in the sources, with $|{S}\rangle=|\tilde{{S}}\rangle=V|P_{rel},\ell\rangle$ and $P_{rel}\equiv\sqrt{M_{N}E_{rel}}$. Thus, the interaction operators in the $H$ and sources are fixed.
@@
Specifically, when setting up training calculations, we fix $\mathrm{Im}\,E^{tr}_{\alpha}=10$ MeV and sample $N_{b}=20$ training points using Latin Hypercube sampling (LHS) Tang (1993) in the two-dimensional parameter space for $(\mathrm{Re}\,E,E_{rel})$ with $\mathrm{Re}\,E\in[-50,50]$ MeV and $E_{rel}\in[0,100]$ MeV. We construct the emulator and perform emulation at a randomly chosen point corresponding to on-shell scattering: $E=E_{rel}=21$ MeV.
@@
The emulation-related wave functions in the $s$-wave scattering are plotted in Fig. 2. The top panel shows the training solution wave functions, which are bound-state-like states. The bottom panel compares the emulated wave function and the exact solution: the agreement extends well beyond the range of the sources—recall $\Lambda_{2}=200$ MeV.
@@
The emulated wave function eventually approaches zero with $r\to\infty$ since the training wave functions have this asymptotic. However, as explained in Sec. III, the incorrect asymptotics in wave function emulation pose no issue since the $\mathcal{A}$ is computed via the overlaps in Eq. (5). The spatial coverage of the sources is on the order of the interaction range, in which the wave function emulation is correct, and thus, so is the $\mathcal{A}$ emulation.
@@
Finally, the treatment of the $E$ and $E_{rel}$ variables is worth pointing out again, as it is relevant for later discussions. An emulator for varying only $E$ while fixing all the other parameters can provide, for example, $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ along the real axis. Hence, it is suitable for emulating response function calculations. However, for the on-shell scattering amplitudes, if we need to emulate in $E_{rel}$, the $E$ and $E_{rel}$ variables must be varied independently. Otherwise, while keeping the $E=E_{rel}$ condition in the training calculations, the sources are not spatially compact if $E_{rel}$ turns complex-valued.

### v2

@@ III.2 Compressed spectra
[…]
@@
When three conditions are satisfied: $|{S}_{\alpha}\rangle=|\tilde{{S}}_{\alpha}\rangle$ and being invariant under time-reversal $\mathbb{T}$ transformation, and $H$ being Hermitian in the training calculations, $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ become complex symmetric matrices. This property is the shared (but not a required) characteristic in the existing NHQM methods [12, 57]. The training equations turn into,
$(E_{\alpha}-H_{\alpha})|\Psi_{\alpha}^{\mathrm{tr}}\rangle$ $=|{S}_{\alpha}\rangle\ ,$ (31)
$(E_{\alpha}^{\ast}-H_{\alpha})|\tilde{\Psi}_{\alpha}^{\mathrm{tr}}\rangle$ $=|{S}_{\alpha}\rangle\ .$ (32)
Since $\mathbb{T}|S_{\alpha}\rangle=|S_{\alpha}\rangle$, I have
$|\tilde{\Psi}_{\alpha}^{\mathrm{tr}}\rangle$ $=\mathbb{T}|\Psi_{\alpha}^{\mathrm{tr}}\rangle\ ,$ (33)
$|{\Psi}_{\alpha}^{\mathrm{tr}}\rangle$ $=\mathbb{T}|\tilde{\Psi}_{\alpha}^{\mathrm{tr}}\rangle\ .$ (34)
It is then easy to show that
$\begin{bmatrix}N\end{bmatrix}_{\alpha\beta}=\langle\mathbb{T}\Psi_{\alpha}^{\mathrm{tr}}|\Psi_{\beta}^{\mathrm{tr}}\rangle=\left(\langle\Psi_{\alpha}^{\mathrm{tr}}|\mathbb{T}^{\dagger}\Psi_{\beta}^{\mathrm{tr}}\rangle\right)^{\ast}=\begin{bmatrix}N\end{bmatrix}_{\beta\alpha}$ (35)
and similarly $\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}=\begin{bmatrix}H\end{bmatrix}_{\beta\alpha}$. This proves the previous assertion. Note that $H$ is Hermitian in the training calculations, but $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ are non-Hermitian in general and complex symmetric in this particular case.
As briefly mentioned at the beginning of Sec. III and elaborated later in Secs. IV and V, these eigenvalues are distributed in a pattern similar to the pole distribution of a (near-)optimal rational approximation discussed in Sec. II.3. The emulation performance certainly depends on whether $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ can reproduce the physics, including the pole positions, in the region where the bound states and resonances are located.
However, the DC states in the emulator’s compressed spectrum depend on the setup of the training points, in particular, the distribution of $E^{\mathrm{tr}}_{\alpha}$ in the complex $E$ plane. As one interesting specific case, when (1) $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ are complex symmetric and (2) both $E^{\mathrm{tr}}_{\alpha}$ and $E^{\mathrm{tr}\ast}_{\alpha}$ are included as the training points associated with a given $\bm{\theta}_{\alpha}$, the eigenenergies are always real and thus the discretized BCTs are along the real axis. See Fig. 3 for a visual presentation, which is a special case with $\bm{\theta}$ fixed. This point is further elaborated in the Appendix A. Such a setup could be beneficial for discrete bound spectrum calculations. However, the discretized BCTs should be away from the real axis for the continuum calculations. Therefore, I focus on the emulations with $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}>0$ in the training points.
@@ III.3 Difference between the existing NHQM and the CEE methods
I am now at a good point to discuss the differences between the existing NHQM methods and the CEE. The previous approaches must analytically continue the conventional single-particle basis (e.g., harmonic oscillator basis) in which $H$ is Hermitian to a basis that $H$ turns non-Hermitian. These methods then construct a many-body basis as a direct product of the single-particle basis. The numerical difficulty could arise during the analytical continuation, limiting how far away from the real axis the BCTs can be moved.
In contrast, CEE achieves the “non-Hermitization” of $H$ by building the many-body basis from the solution snapshots of the inhomogeneous Schrödinger equations with complex energies, i.e., by the RBM-based subspace projection (or viewed as spectrum compression) during the on-line emulation stage. Note these snapshot basis build in inter-particle correlations.
At the off-line training stage, one does not need to—but one can—perform analytical continuation of the $H$ matrix to obtain the training solution snapshots, as the solutions to the equation with a Hermitian $H$ already have bound-state-like boundary conditions. As a result, the non-Hermitian $\begin{bmatrix}H\end{bmatrix}$ matrices presented in this work generally have much lower dimensions than the corresponding matrix in the other NHQM methods.
It is also relevant to note that my approach targets a specific portion of the spectrum, as the eigenenergies of the compressed spectrum are distributed in a finite domain of the complex $E$ plane. The portion is not precisely defined. However, when the $E$ variable is extrapolated too far away from the range of the training energies during the emulation stage, the emulation errors become significant. In contrast, in the existing NHQM methods, the spectrum range is defined by the energy cut-off applied on a single-particle basis.
@@ III.4 Some numerical details
At the training stages, I typically over-sampled training points in the parameter space for simplicity. I.e., $N_{b}$ is larger than the true dimension of the solution subspace. As the result, during emulations, the system of Eq. (19) becomes ill-conditioned. This is not a deficiency of the framework but, in fact, tied to the efficiency of the dimension reduction. However, regularization needs to be implemented to remove the redundancy. Here, I apply the singular value decomposition (SVD) to the $\begin{bmatrix}M\end{bmatrix}$ and eliminate the singular values $\sigma_{.}$ that satisfy $|\sigma_{i}/\sigma_{max}|<10^{-13}$, a threshold similar to the relative errors of the training calculation, to perform pseudo-inverse to compute $c$ and $\tilde{c}$. The truncation on the singular values leads to an effective number of poles in the emulated $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ in the complex-$E$ plane, which is labeled as $N_{\mathrm{eff}}$.
For the same reason, when solving the generalized eigenvalue problem to emulate spectra using Eq. (30), I invert $\begin{bmatrix}N\end{bmatrix}$ also using the pseudo-inverse with the same SVD singular value truncation, which again reveal the dimension of the subspace ($N_{\mathrm{eff}}$). Bear in mind that it is only when the training points are over-sampled that $N_{\mathrm{eff}}$ approaches the dimension of the subspace; otherwise, $N_{\mathrm{eff}}$ is the same as $N_{b}$.
In the future, the so-called greedy algorithms [107, 20, 108] should be pursued to strategically and efficiently sample the training points, which stops at the right $N_{b}$ that is close to the dimension of the solution subspace.
It is worth pointing out a second way to emulate $\mathcal{A}$ in $E$ when one has an emulated spectrum at $\bm{\theta}$. Note that
@@
with $E_{\beta}$ as the emulated eigenvalues. To fix all the residues, $W_{\beta}$, I first compute $\mathcal{A}(E_{\alpha}^{\mathrm{tr}},\bm{\theta})$ by using the first $\mathcal{A}$ emulation method (i.e., solving Eq. (19) and then evaluating Eq. (24)). The residues can be computed by solving a linear equation system:
$\begin{bmatrix}\mathcal{A}(E_{\alpha}^{\mathrm{tr}},\bm{\theta})\end{bmatrix}$ $=\sum_{\beta=1}^{N_{b}}\frac{W_{\beta}(\bm{\theta})}{E_{\alpha}^{\mathrm{tr}}-E_{\beta}(\bm{\theta})}\ .$ (37)
@@ III.5 Existing works
In the MOR studies of dynamic linear systems, the equations in the frequency domain take the form of Eq. (10). The so-called rational-Krylov method (e.g., Chap. 10 and 11 in Ref. [46]) is similar to the CEE. However, these studies are about systems with finite-dimensional matrices. Similarly, recent two studies [47, 48] in quantum chemistry also apply this method to emulate $\mathcal{A}(E)$ involving continuum physics. Again, the critical difference between finite-dimensional matrix and $H$ with continuum was not studied there, including the point that the continuum spectra can be efficiently compressed. It seems the emulated eigenenergies in these works are always real, i.e., their Galerkin projection, used to derive the ROMs, differs significantly from the one used in this work. In addition, I demonstrate the effectiveness of emulation for other parameters, which has not been studied in Refs. [47, 48].
@@ IV Two-body demonstrations
[…]
@@
The form factor in the momentum-space representation, $g_{\ell}(q)$, is a Gaussian function with a width parameter $\Lambda_{2}$; $\lambda$ is the coupling-strength parameter. Here, I focus on $s$ and $p$ waves. The associated $\mathcal{A}(E)$s are known analytically in the complex $E$ plane. The details about the model and relevant analytical formulas can be found in Appendix B.
For emulation, the training calculations are carried out using a bound-state method, employing the Lagrange function of the Legendre polynomials as bases [80]. The basis function is defined in a finite interval of the spatial radial coordinate, $0,R_{IR}]$, as mentioned in discussing Fig. 1. I choose $R_{IR}$ large enough and the UV resolution scale small enough (via increasing the basis size) so that the emulator components, $\begin{bmatrix}N\end{bmatrix}$ and $\begin{bmatrix}H\end{bmatrix}$, all converged to a level of $10^{-12}$ or even smaller.
@@ IV.1 Emulation with fixed $H$ but varying sources
Figure 2: Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{\mathrm{rel}})$. The training wave functions are plotted in the top panel, and at the bottom, the emulated and the exact wave functions are compared for a particular point for on-shell scattering. Note that all the plotted function are $\langle r|\Psi\rangle$ multiplied by $r$. The information about training points is listed in the top panel, while the information about the testing point is provided in the title of the bottom panel.
@@
In this subsection, I hope to gain some understanding of my emulators by analyzing the basic behaviors of the wave function solution of Eq. (2) in coordinate space. Here, the sources are chosen as $|{S}\rangle=|\tilde{{S}}\rangle=V|P_{\mathrm{rel}},\ell\rangle$ with $P_{\mathrm{rel}}\equiv\sqrt{M_{N}E_{\mathrm{rel}}}$. Bear in mind that $E_{\mathrm{rel}}$ is always positive.
To begin with, some analytical analysis is in order here. In my simple example, the sources can be expressed as
$|S\rangle$ $=|g\rangle\times\left(\lambda\langle g|P_{\mathrm{rel}},\ell\rangle\right)\ ,$ (39)
which are localized, as long as the overlap integral in the expression is finite—as guaranteed by keeping $E_{\mathrm{rel}}$ and thus $P_{\mathrm{rel}}$ real. To deal with the exact solution $|\Psi(E)\rangle$, it is rewritten as
$|\Psi(E)\rangle$ $=\frac{1}{E-H_{0}}\left(|{S}\rangle+T(E)\frac{1}{E-H_{0}}|{S}\rangle\right)$
[…]
$\equiv\frac{1}{E-H_{0}}|\overline{{S}}\rangle\ .$ (40)
In the second step, both $\tau(E)$ (discussed in Appendix B) and the matrix element of $H_{0}$’s resolvent are smooth functions of $E$ when $E$ is away from the real-$E$ axis. The compactness of $|\overline{{S}}\rangle$ then follows from that of $|g\rangle$. Therefore, at large $r$, based on the asymptotic behavior of the free Green’s function [5], one can see that for $s$-wave,
$\langle r|\Psi(E)\rangle$ $\xrightarrow{r\to\infty}\frac{\exp\left(ipr\right)}{r}F(E)\ ,$ (41)
with the factor $F$ determined by $|\overline{S}\rangle$. Here, $p=\sqrt{2\mu|E|}\exp(i\theta_{E}/2)$ and $\theta_{E}\equiv\arg(E)\in[0,2\pi)$ (i.e, with the BCT defined on the positive-$E$ axis). When $E$ is right above the BCT, $\langle r|\Psi(E)\rangle$ behaves as an outgoing wave. However, if $\arg(E)\neq 0$, the function is exponentially damped at large $r$ due to the $\exp(-r\mathrm{Im}\,p)$ factor.
@@
Now, let me discuss numerical tests on these analytical understandings, for which I consider the $s$-wave channel and choose $\Lambda_{2}=200$ MeV and $\lambda$ such that the system has a bound ground state with a $10$ MeV binding energy. The emulation is performed in a three-dimensional parameter space for $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{\mathrm{rel}})$, while the interaction operators in the $H$ and sources are fixed. Specifically, when setting up training calculations, I fix $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=10$ MeV and sample $N_{b}=20$ training points using Latin Hypercube sampling (LHS) [109] in the two-dimensional parameter space for $(\mathrm{Re}\,E,E_{\mathrm{rel}})$ with $\mathrm{Re}\,E\in[-50,50]$ MeV and $E_{\mathrm{rel}}\in[0,100]$ MeV. I then construct the emulator and perform emulation at a randomly chosen point corresponding to on-shell scattering: $E=E_{\mathrm{rel}}=21$ MeV.
@@
The emulation-related wave functions in the $s$-wave scattering are plotted in Fig. 2 (an extra $r$ factor is included in the plotted functions). The top panel shows the training solutions, which are spatially localized (bound-state-like) states. The bottom panel compares the emulated wave function and the exact solution: the agreement extends well beyond the range of the sources—recall $\Lambda_{2}=200$ MeV. More results with other $E_{\mathrm{rel}}$ values can be found in the Supplemental Material (SM) [110].
@@
The emulated solution eventually approaches zero with $r\to\infty$ since the training solutions have this asymptotic. However, as explained in Sec. III, the incorrect asymptotics in wave function emulation pose no issue since the $\mathcal{A}$ is computed via the overlaps in Eq. (4). The spatial coverage of the sources is on the order of the interaction range, in which the wave function emulation is correct, and thus, so is the $\mathcal{A}$ emulation.
Although these results, including Eqs. (39)-(41), are based on a separable interaction, the asymptotic behaviors of the training and emulated wave functions, as observed here, should be applicable to non-separable interactions as well. To illustrate this point, the so-called Minnesota potential [111] is studied in the SM [110].
@@
Finally, the treatment of the $E$ and $E_{\mathrm{rel}}$ variables is worth pointing out again, as it is relevant for later discussions. An emulator for varying only $E$ while fixing all the other parameters can provide, for example, $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ along the real axis. Hence, it is suitable for emulating response function calculations. However, for the on-shell scattering amplitudes, if one needs to emulate in $E_{\mathrm{rel}}$, the $E$ and $E_{\mathrm{rel}}$ variables must be varied independently. Otherwise, while keeping the $E=E_{\mathrm{rel}}$ condition in the training calculations, the sources are not spatially compact if $E_{\mathrm{rel}}$ turns complex-valued.